# Deep Discrete Encoders: Identifiable Deep Generative Models for Rich Data with Discrete Latent Layers

### v1

@@ 3.1 Model Identifiability
    After permuting the rows, we can write $\mathbf{G}^{(d)\star}=\left[\mathbf{G}_{1}^{(d)\top},\mathbf{G}_{2}^{(d)\top},\mathbf{G}_{3}^{(d)\top}\right]^{\top}$, where $\mathbf{G}_{1}^{(d)}$ and $\mathbf{G}_{2}^{(d)}$ are $K^{(d)}\times K^{(d)}$ matrices whose diagonal entries are ones:
    $\mathbf{G}_{1}^{(d)},~{}\mathbf{G}_{2}^{(d)}=$ $\begin{pmatrix}1&*&\cdots&*\\ *&1&\cdots&*\\ \vdots&\vdots&\ddots&\vdots\\ *&*&\cdots&1\end{pmatrix}$, where each * denotes an arbitrary value in $\{0,1\}$, and $\mathbf{G}_{3}^{(d)}$ does not have any all-zero columns.
@@
Condition C relaxes conditions A and B in Theorem 1. While condition A requires two exclusive children for each latent variable, condition C does not require any exclusive child and allows more complicated dependence structures. Additionally, condition B for the remaining $J-2K^{(d)}$ variables is relaxed to a simple non-zero column condition on $\mathbf{G}^{(d)}_{3}$. Thus, condition C does not concern any continuous parameter values and just depends on the graph structure $\mathbf{G}^{(d)}$, and can be used as a practical criterion to assess identifiability.
@@
We place our identifiability results in the literature on the identifiability of generative models and latent variable models. While the identifiability of generative models has attracted increasing attention in machine learning (Hyvarinen et al.,, 2019; Khemakhem et al.,, 2020; Moran et al.,, 2022; Kivva et al.,, 2022), many of these results require additional information such as auxiliary covariates. More importantly, almost all of these results build on nonlinear independent component analysis (Oja and Hyvarinen,, 2000) or variational autoencoders (Ranzato and Szummer,, 2008), both of which essentially have only one latent layer of random variables transformed by deterministic deep neural networks. Consequently, these results cannot be applied to DDEs with multiple latent layers organized in a probabilistic graphical model. Since uncertainty occurs and accumulates in each layer of a DDE, addressing identifiability in such cases requires fundamentally different techniques due to a complicated marginal likelihood. At the high level, our proof techniques are based on transforming the marginal distribution of data into a tensor and invoking the uniqueness of tensor decompositions to establish identifiability. In statistics, the study of identifiability has a long history but also mainly concerns relatively simple latent structures with only one latent layer (Anderson and Rubin,, 1956; Koopmans and Reiersol,, 1950; Allman et al.,, 2009; Xu and Shang,, 2018). It has been largely unknown whether DGMs with many nonlinear latent layers can be identifiable for general data as considered in this paper.
@@ 3.2 Estimation Consistency
where $\ell(\boldsymbol{\Theta}\mid\mathbf{Y})=\sum_{i=1}^{N}\log\mathbb{P}(\mathbf{Y}_{i}\mid\boldsymbol{\Theta})$ denotes the marginal log-likelihood function given a sample $\mathbf{Y}_{1},\ldots,\mathbf{Y}_{N}$ of size $N$, defined based on the marginal distribution of $\mathbf{Y}$ in (5). From now on, we slightly abuse notation and let $\mathbf{Y}$ denote the $N\times J$ data matrix including $\mathbf{Y}_{1},\ldots,\mathbf{Y}_{N}$ as rows. Using the estimated coefficients in ^$\widehat{\mathcal{B}}=(\widehat{\beta}_{l,k}^{(d)})$, the layer-wise graphical matrices in $\mathcal{G}$ can be estimated by reading off the sparsity pattern of $\widehat{\mathcal{B}}$:
@@
For some tuning parameters $\lambda_{N},\tau_{N}>0$, $p_{\lambda_{N},\tau_{N}}:\mathbb{R}\to[0,\infty)$ is a sparsity-inducing symmetric penalty that is nondecreasing on $[0,\infty)$, nondifferentiable at 00, differentiable at $(0,\tau_{N})$, $p_{\lambda_{N},\tau_{N}}\propto{\lambda_{N}}/{\tau_{N}}$ around 00 and satisfy
Note that $\lambda_{N}$ is the magnitude of the penalty, and $\tau_{N}$ is the point of truncation. Our assumption for the penalty $p_{\lambda_{N},\tau_{N}}$ includes common truncated sparsity-inducing penalties such as the TLP (Truncated Lasso Penalty; Shen et al.,, 2012) and SCAD (Smoothly Clipped Absolute Deviation; Fan and Li,, 2001). With a slight abuse of notation, in (6), we view the penalty $p_{\lambda_{N},\tau_{N}}$ as a function of matrices by letting it be the sum of the entrywise penalties: $p_{\lambda_{N},\tau_{N}}(\mathbf{B}^{(d)})=\sum_{k\in[K^{(d-1)}],\ell\in[K^{(d)}]}p_{\lambda_{N},\tau_{N}}(\beta_{k,\ell}^{(d)}).$
@@
[…]
Remark 1.
One natural question is to whether our estimator would still be consistent when the number of latent variables are unknown. This extension is not straightforward since the number of the top-layer latent variable, $K^{(D)}$, determines the number of deepest mixture components of DDEs. Estimating the number of mixture components is a challenging problem even in simple parametric models, and often leads to a slower (than ${1}/{\sqrt{N}}$) rate of convergence in parameter estimation (Goffinet et al.,, 1992; Ho and Nguyen,, 2016). Note that for such cases, the Fisher information becomes singular, and Theorem 3 cannot be applied. We will give a practical method to choose $\mathcal{K}$ at the end of Section 4.
@@ 4 Scalable Computation Pipeline
4.1 Warm-up Example: Penalized EM Algorithm
This subsection first presents a penalized EM algorithm as a warm-up example, before presenting the scalable computational pipelines in Sections 4.2 (stage one) and 4.3 (stage two).
We describe the methods for two-layer DDEs; they are conceptually easy to generalize to DDEs with more layers. Since the coefficients $\mathbf{B}^{(1)}$ and $\mathbf{B}^{(2)}$ are potentially sparse, we propose a penalized EM algorithm (PEM) to compute the penalized maximum likelihood estimator in (6). The PEM is an iterative procedure that consists of an expectation step followed by a penalized maximization step (Green,, 1990; Chen et al.,, 2015). In the $(t+1)$th iteration, the E-step computes the expectation of the complete data penalized-log-likelihood
${\boldsymbol{p}}^{[t+1]}:=$ $\operatorname*{argmax}_{{\boldsymbol{p}}}\sum_{i=1}^{N}\mathbb{E}\left[\log\mathbb{P}(\mathbf{A}_{i}^{(2)};{\boldsymbol{p}});{\boldsymbol{p}}^{[t]}\right],$ (8)
$\mathbf{B}^{(2),[t+1]}:=$ $\operatorname*{argmax}_{\mathbf{B}^{(2)}}\sum_{i=1}^{N}\mathbb{E}\left[\log\mathbb{P}(\mathbf{A}_{i}^{(1)}\mid\mathbf{A}_{i}^{(2)};\mathbf{B}^{(2)});\mathbf{B}^{(2),[t]}\right]-p_{\lambda_{N},\tau_{N}}(\mathbf{B}^{(2)}),$ (9)
$(\mathbf{B}^{(1),[t+1]},\boldsymbol{\gamma}^{[t+1]}):=$ $\operatorname*{argmax}_{\mathbf{B}^{(1)},\boldsymbol{\gamma}}\sum_{i=1}^{N}\mathbb{E}\left[\log\mathbb{P}(\mathbf{Y}_{i}\mid\mathbf{A}_{i}^{(1)},\mathbf{A}_{i}^{(2)});\mathbf{B}^{(1),[t]},\boldsymbol{\gamma}^{[t]})\right]-p_{\lambda_{N},\tau_{N}}(\mathbf{B}^{(1)}).$ (10)
Here, the optimizations for each layer are separated; we update the top-layer proportion parameters ${\boldsymbol{p}}$ in (8), the middle latent layer coefficients $\mathbf{B}^{(2)}$ in (9), and the bottom layer coefficients $(\mathbf{B}^{(1)},\boldsymbol{\gamma})$ in (10). Additionally, due to the conditional independence assumption in each layer, the maximizations can be further simplified into low-dimensional optimizations over each row of $\mathbf{B}$. Algorithm 1 summarizes the PEM algorithm. The details of the simplified M-step and the implementation are in Supplementary Material S.3.1 and S.3.3.
Data: $\mathbf{Y},\mathcal{K}$, tuning parameters $\lambda_{N},\tau_{N}$.
while log-likelihood has not converged do
       In the $[t+1]$th iteration,
      // E-step
       end for
      // M-step
       update $\boldsymbol{\Theta}^{[t+1]}=({\boldsymbol{p}}^{[t+1]},\mathbf{B}^{(1),[t+1]},\mathbf{B}^{(2),[t+1]},\boldsymbol{\gamma}^{[t+1]})$ by solving (8)-(10).
end while
@@ 4.2 Stage One: Layerwise Double-SVD Initialization
The multiple layers of nonlinearity in DDEs lead to a highly nonconvex optimization landscape with potentially exponentially many local optima (Sutskever et al.,, 2013). If the penalized EM algorithm starts with an initialization close to a local optima, it can get stuck and fail to converge to the global maximizer of the penalized log-likelihood function. Hence, for highly complex latent variable models such as DDEs, it is critical to initialize the optimization algorithm wisely. We propose a novel layerwise nonlinear spectral initialization strategy, which enjoys low computational complexity and reasonably high accuracy. This spectral initialization serves as the first stage in the proposed computational pipeline.
@@ 4.3 Stage Two: Penalized SAEM Algorithm
For DDEs with a large number of latent variables, the E-step in Algorithm 1 requires computing all conditional probabilities $\mathbb{P}(\mathbf{A}^{(1)}_{i}=\boldsymbol{\alpha}^{(1)},\mathbf{A}^{(2)}_{i}=\boldsymbol{\alpha}^{(2)}\mid\mathbf{Y};\boldsymbol{\Theta}^{[t]})$ for all $\boldsymbol{\alpha}^{(1)}\in\{0,1\}^{K^{(1)}}$ and $\boldsymbol{\alpha}^{(2)}\in\{0,1\}^{K^{(2)}}$. This requires computing and storing $O(N\times 2^{\sum_{d=1}^{D}K^{(d)}})$ terms. The exponential dependency in $K^{(d)}$ is concerning even for moderately large latent dimensions, say $K^{(d)}=10$, and quickly becomes prohibitive for larger $K^{(d)}$. Therefore, we propose a penalized Stochastic Approximate EM (SAEM; see Delyon et al.,, 1999; Kuhn and Lavielle,, 2004) by modifying both the E-step and M-step to more scalable versions using approximate sampling. As we illustrate below, this is a method with linear dependence of $\sum_{d=1}^{D}K^{(d)}$ on the computing time as well as memory, and is a scalable alternative to Algorithm 1.
We elaborate on the details on how to modify the PEM to derive the SAEM. First, we replace the E-step to a simulation step, which consists of simulating only a small number (denoted as $C$) of posterior samples of the latent variables. As exact sampling from the joint distribution $\mathbb{P}(\mathbf{A}_{i}^{(1)},\mathbf{A}_{i}^{(2)}\mid\mathbf{Y};\boldsymbol{\Theta}^{[t]})$ is expensive, we sample each latent variable from their complete conditionals. That is, we sample each $A_{i,k}^{(1),[t+1]}$ from $\mathbb{P}(A_{i,k}^{(1)}\mid(-),\boldsymbol{\Theta}^{[t]})$, where ()(-)( - ) denotes the estimates of all other latent variables from the $t$th iteration. Since the latent variables are binary, the conditional distribution is Bernoulli and easy to evaluate. Consequently, the computationally expensive E-step is replaced by the simulation step, which computes and stores only $O(N\times\sum_{d=1}^{D}K^{(d)})$ terms. In terms of choosing the number of samples $C$ in each iteration, we empirically find that taking $C=1$ is computationally efficient without sacrificing much accuracy. Additional simulation studies in Supplementary Material S.4.2 demonstrate that taking a larger $C$ results in only little improvement in accuracy, but becomes significantly slower. This choice of $C=1$ was also suggested in the original paper that proposed the SAEM (Delyon et al.,, 1999).
Second, we modify the M-step objective function by replacing the exact conditional probability values (denoted as $\varphi^{[t+1]}$ in Algorithm 1) to sample-based quantities, and also stochastically averaging the objective functions. For example, we modify (9) as follows:
@@
      for $(i,k)\in[N]\times[K^{(1)}]$ do
             Sample each $A_{i,k}^{(1),[t+1]}$ from the complete conditional $\mathbb{P}(A_{i,k}^{(1)}\mid(-))$ using the previous parameter estimates $\boldsymbol{\Theta}^{[t]},\mathbf{A}^{[t]}$
      for $(i,l)\in[N]\times[K^{(2)}]$ do
             Sample each $A_{i,l}^{(2),[t+1]}$ from the complete conditional $\mathbb{P}(A_{i,l}^{(2)}\mid(-))$ using the previous parameter estimates $\boldsymbol{\Theta}^{[t]},\mathbf{A}^{[t]}$
@@
We similarly modify the optimizations in (8)–(10). Algorithm 3 summarizes our final SAEM algorithm with $C=1$, where exact formulas are deferred to Supplement S.3.1. Similar to the PEM algorithm, the M-step updates can be written in terms of low-dimensional maximizations for each row of the coefficient matrices.
Selecting the latent dimension $\mathcal{K}$.
To apply the above computation pipeline to real data, one additionally needs to specify the number of latent variables, $\mathcal{K}$. For this purpose, we propose a layer-wise estimation strategy, which can be incorporated into our initialization procedure in Algorithm 2. Recall the denoised data matrix $\widehat{\mathbf{Z}}$ from Step 1 of Algorithm 2 and let $\sigma_{1},\sigma_{2},\ldots$ be its singular values in the descending order. Given a candidate grid $\mathfrak{K}$, we estimate the number of latent variables in the first layer based on the largest spectral ratio: $\widehat{K}^{(1)}:=\operatorname*{argmax}_{k\in\mathfrak{K}}({\sigma_{k}}/{\sigma_{k+1}})-1$. Now, given $\widehat{K}^{(1)}$, we proceed with the remaining steps of Algorithm 2 to estimate the first-layer latent variables $\mathbf{A}^{(1)}$. By treating the estimated $\mathbf{A}^{(1)}$ as the observed variables of a one-latent-layer DDE, we can repeat the above procedure to estimate $\widehat{K}^{(2)}$. See Supplement S.3.2 for alternative selection criteria for $\mathcal{K}$.

### v2

@@ 3.1 Model Identifiability
    There exists a partition of the (d1)(d-1)th layer variable indices $K^{(d-1)}]=\mathcal{I}_{1}\cup\mathcal{I}_{2}\cup\mathcal{I}_{3}$ that satisfies the following properties: (i) for $a=1,2$, there is a perfect matching in the bipartite graph between $\mathbf{A}^{(d-1)}_{\mathcal{I}_{a}}$ and $\mathbf{A}^{(d)}$, (ii) each latent variable $A_{k}^{(d)}$ has at least one child among $\mathbf{A}^{(d-1)}_{\mathcal{I}_{3}}$.
@@
Condition C relaxes conditions A and B in Theorem 1. While condition A requires two pure children for each latent variable, condition C does not require any pure child and allows more complex dependence structures. Additionally, condition B for the remaining $J-2K^{(d)}$ variables is relaxed to a simple non-zero column condition on $\mathbf{G}^{(d)}_{3}$. Thus, condition C does not concern any continuous parameter values and just depends on the graph structure $\mathbf{G}^{(d)}$, and can be used as a practical criterion to assess identifiability.
Example 1 (Illustration of identifiability conditions).
We provide a toy example to illustrate and compare conditions in Theorem 1 and Theorem 2. In Figure 3, condition A holds when only the solid edges exist; in this case, the latent variable $A_{1}^{(d)}$ has two pure children $A_{1}^{(d-1)},A_{3}^{(d-1)}$, and latent variable $A_{2}^{(d)}$ also has two pure children. In contrast, condition C allows arbitrary additional dashed arrows in the graph structure; in this case, by taking $\mathcal{I}_{1}=\{A_{1}^{(d-1)},A_{2}^{(d-1)}\}$ and $\mathcal{I}_{2}=\{A_{3}^{(d-1)},A_{4}^{(d-1)}\}$, the red and blue solid edges each form a perfect matching. Thus, condition C significantly relaxes the pure child condition in condition A, by allowing many more additional edges in the bipartite graph.
$A_{1}^{(d-1)}$$A_{2}^{(d-1)}$$A_{3}^{(d-1)}$$A_{4}^{(d-1)}$$A_{5}^{(d-1)}$$A_{1}^{(d)}$$A_{2}^{(d)}$
[…]
Figure 3: Graphical illustration of conditions A and C. While condition A holds when all dashed arrows are ignored (or equivalently, =0*=0 in the matrix representation), condition C can allow arbitrarily many dashed arrows (where each * can be either zero or one).
@@
We place our identifiability results in the literature on the identifiability of generative models and latent variable models. While the identifiability of generative models has attracted increasing attention in machine learning (Hyvarinen et al., , 2019; Khemakhem et al., , 2020; Moran et al., , 2022; Kivva et al., , 2022), many of these results require additional information such as auxiliary covariates. More importantly, almost all of these results build on nonlinear independent component analysis (Oja and Hyvarinen, , 2000) or variational autoencoders (Ranzato and Szummer, , 2008), both of which essentially have only one latent layer of random variables transformed by deterministic deep neural networks. Consequently, these results cannot be applied to DDEs with multiple latent layers organized in a probabilistic graphical model. Since uncertainty occurs and accumulates in each layer of a DDE, addressing identifiability in such cases requires fundamentally different techniques due to a complicated marginal likelihood. At the high level, our proof techniques are based on transforming the marginal distribution of data into a tensor and invoking the uniqueness of tensor decompositions to establish identifiability.
In statistics, the study of identifiability has a long history but also mainly concerns relatively simple latent structures with only one latent layer (Anderson and Rubin, , 1956; Koopmans and Reiersol, , 1950; Allman et al., , 2009; Xu and Shang, , 2018). In particular, Gu and Dunson, (2023) establish identifiability of a multi-layer latent structure model, but their result is only applicable to categorical response data, in contrast to the general and rich types of responses considered in this work; moreover, Gu and Dunson, (2023)’s strict identifiability result requires each latent variable to have at least three pure children. On the other hand, Lee and Gu, (2024) considered models for general responses which are similar to ours, but their model is restricted to only one latent layer, and they prove identifiability under a strong assumption that the true graph between the latent and observed layers is known. Compared to these existing results for related models, our results (i) apply to arbitrary response types as opposed to only categorical responses, (ii) require weaker identifiability conditions compared to Gu and Dunson, (2023) (in terms of both strict and generic identifiability), and (iii) do not require a known graph structure as in Lee and Gu, (2024).
@@ 3.2 Estimation Consistency
where $\ell(\boldsymbol{\Theta}\mid\mathbf{Y})=\sum_{i=1}^{N}\log\mathbb{P}(\mathbf{Y}_{i}\mid\boldsymbol{\Theta})$ denotes the marginal log-likelihood function given a sample $\mathbf{Y}_{1},\ldots,\mathbf{Y}_{N}$ of size $N$, defined based on the marginal distribution of $\mathbf{Y}$ in (5). From now on, we slightly abuse notation and let $\mathbf{Y}$ denote the $N\times J$ data matrix including $\mathbf{Y}_{1},\ldots,\mathbf{Y}_{N}$ as rows. Using the estimated coefficients in $\widehat{\mathcal{B}}=(\widehat{\beta}_{l,k}^{(d)})$, the layer-wise graphical matrices in $\mathcal{G}$ can be estimated by reading off the sparsity pattern of $\widehat{\mathcal{B}}$:
@@
For some tuning parameters $\lambda_{N},\tau_{N}>0$, $p_{\lambda_{N},\tau_{N}}:\mathbb{R}\to[0,\infty)$ is a sparsity-inducing symmetric penalty that satisfies several technical conditions postponed to Supplement S.1.4. Our assumption for the penalty $p_{\lambda_{N},\tau_{N}}$ includes common truncated sparsity-inducing penalties such as the TLP (Truncated Lasso Penalty; Shen et al., , 2012) and SCAD (Smoothly Clipped Absolute Deviation; Fan and Li, , 2001). With a slight abuse of notation, in (6), we view the penalty $p_{\lambda_{N},\tau_{N}}$ as a function of matrices by letting it be the sum of the entrywise penalties: $p_{\lambda_{N},\tau_{N}}(\mathbf{B}^{(d)})=\sum_{k\in[K^{(d-1)}],\ell\in[K^{(d)}]}p_{\lambda_{N},\tau_{N}}(\beta_{k,\ell}^{(d)}).$
@@
Consider a $D$-latent-layer DDE with true parameters $\boldsymbol{\Theta}^{\star},\mathcal{G}^{\star}$ and known $\mathcal{K}$. Assume that the model at $\boldsymbol{\Theta}^{\star}$ is identifiable, has a non-singular Fisher information, and all entries of $\{\mathbf{B}^{(d)}\}_{d=1}^{D}$ are bounded. Then, the estimator $\widehat{\boldsymbol{\Theta}}$ in (6) is $\sqrt{N}$-consistent in the sense that there exists some $\widetilde{\boldsymbol{\Theta}}\sim_{\mathcal{K}}\widehat{\boldsymbol{\Theta}}$ such that $\lVert\widetilde{\boldsymbol{\Theta}}-\boldsymbol{\Theta}^{\star}\rVert=O_{p}({1}/{\sqrt{N}})$. Here, $\|\cdot\|$ denotes the vectorized $L^{2}$ norm. Additionally, the graphical matrices are consistently estimated: for $\widetilde{\mathcal{G}}$ resulting from $\widetilde{\boldsymbol{\Theta}}$ according to (7), we have $\mathbb{P}(\widetilde{\mathcal{G}}\neq\mathcal{G}^{\star})\rightarrow 0$.
@@ 4 Scalable Computation Pipeline
We present a two-stage scalable computational pipeline to compute the penalized maximum likelihood estimator in (6). Our proposed method builds upon the standard penalized EM algorithm (see Supplement S.3.1) by including a spectral initialization and stochastic approximation. We separate out each stage into Sections 4.1 (stage one) and 4.2 (stage two). For notational simplicity, we describe the proposed method assuming $D=2$ latent layers, which straightforwardly extends to an arbitrary number of latent layers.
@@ 4.1 Stage One: Layerwise Double-SVD Initialization
The multiple layers of nonlinearity in DDEs lead to a highly nonconvex optimization landscape with potentially exponentially many local optima (Sutskever et al., , 2013). For example, if the EM algorithm starts with an initialization close to a local optima, it can get stuck and fail to converge to the global maximizer of the penalized log-likelihood function. Hence, for highly complex latent variable models such as DDEs, it is critical to initialize the optimization algorithm wisely. We propose a novel layerwise nonlinear spectral initialization strategy, which enjoys low computational complexity and reasonably high accuracy. This spectral initialization serves as the first stage in the proposed computational pipeline.
@@ 4.2 Stage Two: Penalized SAEM Algorithm
For DDEs with a large number of latent variables, the E-step in standard EM algorithms (see Supplement S.3.1) computes all conditional probabilities $\mathbb{P}(\mathbf{A}^{(1)}_{i}=\boldsymbol{\alpha}^{(1)},\mathbf{A}^{(2)}_{i}=\boldsymbol{\alpha}^{(2)}\mid\mathbf{Y};\boldsymbol{\Theta}^{[t]})$ for all $\boldsymbol{\alpha}^{(1)}\in\{0,1\}^{K^{(1)}}$ and $\boldsymbol{\alpha}^{(2)}\in\{0,1\}^{K^{(2)}}$. This requires computing and storing $O(N\times 2^{\sum_{d=1}^{D}K^{(d)}})$ terms. The exponential dependency in $K^{(d)}$ is concerning even for moderately large latent dimensions, say $K^{(d)}=10$, and quickly becomes prohibitive for larger $K^{(d)}$. Therefore, we propose a penalized Stochastic Approximate EM (SAEM; see Delyon et al., , 1999; Kuhn and Lavielle, , 2004) by modifying both the E-step and M-step to more scalable versions using approximate sampling. As we illustrate below, this is a method with linear dependence of $\sum_{d=1}^{D}K^{(d)}$ on the computing time as well as memory.
We elaborate on the details on deriving the SAEM. First, we replace the E-step to a simulation step, which consists of simulating only a small number (denoted as $C$) of posterior samples of the latent variables. As exact sampling from the joint distribution $\mathbb{P}(\mathbf{A}_{i}^{(1)},\mathbf{A}_{i}^{(2)}\mid\mathbf{Y};\boldsymbol{\Theta}^{[t]})$ is expensive, we sample each latent variable from their complete conditionals. That is, we sample each $A_{i,k}^{(1),[t+1]}$ from $\mathbb{P}(A_{i,k}^{(1)}\mid(-),\boldsymbol{\Theta}^{[t]})$, where ()(-) denotes the estimates of all other latent variables from the $t$th iteration. Since the latent variables are binary, the conditional distribution is Bernoulli and easy to evaluate. Consequently, the computationally expensive E-step is replaced by the simulation step, which computes and stores only $O(N\times\sum_{d=1}^{D}K^{(d)})$ terms. In terms of choosing the number of samples $C$ in each iteration, we empirically find that taking $C=1$ is computationally efficient without sacrificing much accuracy (see Supplement S.4.4). This choice of $C=1$ was also suggested in the original paper that proposed the SAEM (Delyon et al., , 1999).
Second, we modify the standard M-step objective function (expected complete data log-likelihood) by (i) replacing the exact conditional probability values to sample-based quantities, (ii) stochastically averaging the objective functions, and (iii) including the sparsity-inducing penalties in (6). The M-step objective for updating $\mathbf{B}^{(2)}$ is
@@
 Sample each $A_{i,k}^{(1),[t+1]},A_{i,l}^{(2),[t+1]}$ from the complete conditionals using the previous parameter estimates $\boldsymbol{\Theta}^{[t]},\mathbf{A}^{[t]}$
@@
Data: $\widehat{\mathbf{A}}^{(0)}:=\mathbf{Y}_{N\times J},\widehat{K}^{(0)}:=J$, the number of latent layers $D$
For each $1\leq d\leq D$, repeat:
  1. 1.
    Let $\mathfrak{K}^{(d)}:=\{\lceil\widehat{K}^{(d-1)}/4\rceil,\ldots,\lfloor\widehat{K}^{(d-1)}/2\rfloor\}$ be the candidate grid for $\widehat{K}^{(d)}$.
  2. 2.
    Define $\widehat{K}^{(d)}$ based on the largest spectral ratio of the denoised/linearized $\widehat{\mathbf{A}}^{(d-1)}$ (see step 1 in Algorithm 1):
    $\widehat{K}^{(d)}:=\operatorname*{argmax}_{k\in\mathfrak{K}^{(d)}}\sigma_{k}/\sigma_{k+1}-1.$
  3. 3.
    Given $\widehat{K}^{(d)}$, estimate the $d$-th layer latent variables $\widehat{\mathbf{A}}^{(d)}_{N\times\widehat{K}^{(d)}}$ using Algorithm 1.
Output : Estimated latent dimensions for all layers $\widehat{K}^{(1)},\ldots,\widehat{K}^{(D)}$.
Algorithm 3 Practical methods for selecting the latent dimensions $\mathcal{K}$
Algorithm 2 summarizes our proposed SAEM algorithm with $C=1$, where detailed formulas are deferred to Supplement S.3.2. Here, the detailed M-step updates can be written in terms of low-dimensional maximizations for each row of the coefficient matrices.
Selecting the latent dimensions.
To apply the above computational pipeline to real data, one also needs to specify the number of latent variables, $\mathcal{K}$. We propose a layer-wise estimation strategy in Algorithm 3, which can be incorporated into our initialization procedure in Algorithm 1. Recall the denoised data matrix $\widehat{\mathbf{Z}}$ from Step 1 of Algorithm 1 and let $\sigma_{1},\sigma_{2},\ldots$ be its singular values in the descending order. Given a candidate grid $\mathfrak{K}^{(1)}$, we estimate the size of the first latent layer based on the largest spectral ratio: $\widehat{K}^{(1)}:=\operatorname*{argmax}_{k\in\mathfrak{K}}({\sigma_{k}}/{\sigma_{k+1}})-1$. Now, given $\widehat{K}^{(1)}$, we proceed with the remaining steps of Algorithm 1 to estimate the first-layer latent variables $\mathbf{A}^{(1)}$. Inductively treating the estimated $\widehat{\mathbf{A}}^{(d)}$ as the observed variables of a one-latent-layer DDE, we can repeat the above procedure to estimate $\widehat{K}^{(d+1)}$. See Supplement S.3.3 for alternative selection criteria for $\mathcal{K}$.